OpenStudy (nbarrera):

WILL GET MEDAL!!! Evaluate √x+7 when x = 9. Write the answer in simplified radical form. A. 16 B. 8 C. 10 D. 4

OpenStudy (anonymous):

just use your calculator and put in \[\sqrt{9+7}\] and it gives you your answer

OpenStudy (nbarrera):

Ooh i'l just add √9+7 which = √16 and then simplifies to 8 Thx:)

OpenStudy (anonymous):

the square root of 16 isn't 8

OpenStudy (nbarrera):

I mean 4 cuz 2*8=16 and simplify the square root of 8 and =4

OpenStudy (anonymous):

haha there you go

OpenStudy (anonymous):

but actually the square root of 16 is 4 because when you set 4^2, you get 16. 8*2 does not equal a square root. A square is when you multiply two of the same numbers